OpenStudy (anonymous):

Select the equation of the line that passes through the point (2, -1) and has slope 3 in point-slope form. (y + 1) = 3(x - 2) (y - 1) = 3(x + 2) (x + 2) = 3(y - 1) (x - 2) = 3(y + 1)

OpenStudy (anonymous):

its the first one :D

OpenStudy (anonymous):

Okay, I got this lol. Formula: y-y1=m(x-x1) y-(-1)=3(x-2) So... Answer: y+1=3(x-2)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yepp thats what i said lol i just didnt show the steps xD

OpenStudy (anonymous):

No problem.